Suppose two stars obey1mass luminosity relationship.one star is four times as massive as the other.how do the luminosities of the two stars compare??


1
Expert's answer
2021-01-06T13:53:30-0500

The mass-luminosity relation gives relationship between the mass and the luminosity of a star. It is slightly different for stars with masses in different ranges, but for stars with nearly solar masses we have the luminosity to be proportional to mass in fourth degree: "L \propto M^4."

So for two stars

"\dfrac{L_1}{L_2} = \left( \dfrac{M_1}{M_2}\right) ^4 = 4^4 =256."

So one of the stars is 256 times more luminous than the other.